OpenStudy (anonymous):

. sq. rt of (x- 5) = 3

OpenStudy (anonymous):

sq. rt of (x- 5) = 3 (x-5)=9 x=14

OpenStudy (anonymous):

\[(\sqrt{x-5})^2=3^2\] \[x-5=9\] x=14 check your answer: sqrt(14-5)=3 sqrt(9)=3? 3=3